OpenStudy (anonymous):

Can someone explain to me what exactly I am doing with this problem? My lesson on this doesn't make sense and I can't find anywhere to get help on showing me how to solve this. Determine the zeros of f(x) = x3 – 3x2 – 16x + 48

OpenStudy (anonymous):

zeros means values of x for which f(x)=0 x3-3x2-16x+48=0 x2(x-3)-16(x-3)=0 (x-3)(x2-16)=0 (x-3)(x+4)(x-4)=0 so the zeroes are 3,4 and -4.

OpenStudy (anonymous):

How do you figure out what x is?

OpenStudy (anonymous):

if product of a few no. is zero that means one of the numbers is zero. so either x-3 =0 (which would mean x=3) or x+4 =0 meaning x=-4 or x-4=0 meaning x=4

OpenStudy (anonymous):

How do you get to the 3, 4, -4

OpenStudy (anonymous):

by solving the eq. first we factorise the eq. and then equate each factor to zero because one of them has to be zero